# From Experts to a Generalist: Toward General Whole-Body Control for Humanoid Robots

### v1

@@ Abstract
Achieving general agile whole-body control on humanoid robots remains a major challenge due to diverse motion demands and data conflicts. While existing frameworks excel in training single motion-specific policies, they struggle to generalize across highly varied behaviors due to conflicting control requirements and mismatched data distributions. In this work, we propose BumbleBee (BB), an expert-generalist learning framework that combines motion clustering and sim-to-real adaptation to overcome these challenges. BB first leverages an autoencoder-based clustering method to group behaviorally similar motions using motion features and motion descriptions. Expert policies are then trained within each cluster and refined with real-world data through iterative delta action modeling to bridge the sim-to-real gap. Finally, these experts are distilled into a unified generalist controller that preserves agility and robustness across all motion types. Experiments on two simulations and a real humanoid robot demonstrate that BB achieves state-of-the-art general whole-body control, setting a new benchmark for agile, robust, and generalizable humanoid performance in the real world.
footnotetext: These authors contributed equally to this work.footnotetext: Correspondence to Zongqing Lu $<$zongqing.lu@pku.edu.cn$>$.
@@ 1 Introduction
Humanity has long dreamed of using humanoid robots to assist or replace humans in completing various daily tasks, including grabbing objects (Lenz et al., 2015; Joshi et al., 2020; Miller and Allen, 2004; Du et al., 2021; Shao et al., 2020), carrying goods (Pereira et al., 2002; Dao et al., 2024; Agravante et al., 2019), etc. At present, a single agile whole-body humanoid skill can be trained to a remarkably high level of performance. The ASAP (He et al., 2025) framework first trains motion-specific tracking policies in simulation, then refines them using real-world data through motion-specific dynamic policy learning. This framework enables the robot to effectively bridge the sim-to-real gap and achieve a wide repertoire of agile and expressive behaviors. However, it is hard to directly apply this framework to general motion tracking, as different motions require the robot to focus on different aspects of control. For example, aggressive actions like jumping or fast walking demand precise, high-torque control, while conservative motions prioritize balance and smoothness. These differences lead to mismatched data distributions across motion types. It can lead to conflicting gradients during training, ultimately degrading the performance of the policy.
@@
Intuitively, to mitigate the issues caused by mismatched data distributions, there are two main directions. One is at the model level (Be more powerful): employing more expressive architectures such as Transformers or Diffusion Models (Shridhar et al., 2023; Kim et al., 2021; Brohan et al., 2022; Radosavovic et al., 2024), which are capable of capturing complex motion distributions across diverse skills. The other is at the data level (Decompose the complexity): introducing structure into the training data, such as clustering motions by type or difficulty, balancing data across motion categories, or using curriculum learning strategies to gradually expose the policy to increasingly diverse behaviors (Ji et al., 2024). These approaches aim to reduce interference between conflicting samples and improve the generalization capability of the learned policy. Inspired by the success of mixture of experts (MoE) (Shazeer et al., 2017; Lepikhin et al., 2020) in LLMs, we follow the second path to build a general whole-body control policy.
@@
To this end, we propose BumbleBee (BB), an expert-generalist pipeline that enables a general agile humanoid whole-body controller. Since both tracking and dynamic policies benefit from training on behaviorally similar motions, the dataset is first clustered using an autoencoder (AE) (Hinton and Salakhutdinov, 2006) trained on kinematic features, including additional leg-related features, and motion text descriptions. This design reflects the fact that different leg movements, such as walking, jumping, or crouching, require distinct torque control patterns, making leg dynamics a key factor in categorizing motion. A general tracking policy is initially trained on the full dataset and then fine-tuned within each cluster to obtain specialized expert policies. These experts are deployed on the real robot to collect real-world trajectories. For each cluster, a delta action model is trained to compensate for sim-to-real discrepancies by modeling the difference between simulated and real-world state transitions. The expert policies are further refined using the corresponding delta model in an iterative manner. Finally, a unified generalist is obtained through a combination of expert knowledge distillation.
@@ 2.1 Humanoid Whole-Body Control
Whole-body control (WBC) (Chignoli et al., 2021; Darvish et al., 2019; Dallard et al., 2023; Hutter et al., 2016; Kajita et al., 2001; Moro and Sentis, 2019; Li et al., 2025; Ding et al., 2025) is essential for high DoF humanoid robots to perform coordinated and stable movements in complex tasks. Recent advances in reinforcement learning have enabled robots to acquire diverse whole-body skills, yet training a general WBC policy remains challenging due to hardware constraints and the inherent complexity of the robotic action space.
Previous works address this in different ways: the H2O series (He et al., 2024a, b) incorporates global keypoint positions but suffers from error accumulation in long-horizon tasks; HumanPlus (Fu et al., 2024) uses only joint angles for better real-world adaptability; Hover (He et al., 2024c) introduces observation masking for flexible control mode switching. However, these methods are all trained on the full AMASS (Mahmood et al., 2019) dataset, where diverse motion distributions may cause gradient conflicts and reduce generalization. Exbody2 (Ji et al., 2024) attempts to mitigate this through progressive learning based on task difficulty.
@@ 2.2 Sim-to-Real Transfer
The sim-to-real transfer challenge primarily arises from two factors: (1) inaccuracies in robot modeling, including structural simplifications and parameter errors, and (2) the complex, nonlinear, and time-varying dynamics of real-world environments, e.g., contact events, friction variations, and elastic deformations, that exceed the fidelity of modern physics simulators.
Traditional approaches address this gap through System Identification (SysID) (Åström and Eykhoff, 1971; Kozin and Natke, 1986; Yang et al., 2024; Yu et al., 2018, 2020), which uses real-world observations to calibrate simulator parameters for improved physical realism. Domain Randomization offers an alternative by perturbing environment parameters during training to enhance policy robustness and generalization.
@@ 3 Method
Dataset Curation.
Following prior work (He et al., 2024a), we first retarget SMPL-format (Loper et al., 2015) human motion sequences from the AMASS dataset into robot-specific representations, including global translations and joint axis-angle rotations. Given that AMASS contains a wide range of motions, such as crawling and climbing, we perform an additional data cleaning step to ensure quality. We apply PHC (Luo et al., 2023) to filter the dataset (He et al., 2025), resulting in 8,179 high-quality trajectories for clustering and training.
3.1 AE Clustering
@@
Our method not only relies solely on autoencoders to reconstruct full motion sequences, but also incorporates textual annotations to align motions at the semantic level. This ensures that motions with diverse patterns but shared semantics are positioned closer in the latent space. For example, the act of walking may take the form of linear movement or circular paths. These patterns would be hard to associate with based only on motion alignment. However, things can be different when textual semantic alignment is introduced. Note that the text information for training is from the open-source HumanML3D (Guo et al., 2022) dataset, which offers textual annotations and corresponding frame ranges for most of the sequences in AMASS.
@@
For the encoding process, we adopt a Transformer for motion encoding, following the previous work (Petrovich et al., 2023, 2022). The motion encoder takes as input a motion sequence $M_{\rm full}=\{\mathbf{p}_{t},\mathbf{r}_{t},\dot{\mathbf{r}}_{t},\mathbf{c}_{t},\mathbf{v}^{\text{feet}}_{t}\}_{t=1}^{T}$ and outputs a motion representation $z^{m}$, where $\mathbf{p}_{t}\in\mathbb{R}^{N\times 3}$ represents the 3D positions of all joints, $\mathbf{r}_{t}\in\mathbb{R}^{3}$ is the root translation in 3D space, $\dot{\mathbf{r}}_{t}\in\mathbb{R}^{3}$ is the root velocity, $\mathbf{c}_{t}\in\{0,1\}^{F}$ denotes the binary contact states of $F$ feet, and $\mathbf{v}^{\text{feet}}_{t}\in\mathbb{R}^{F\times 3}$ represents the 3D velocities of the $F$ feet. Textual data $l$ is first processed through a BERT model (Devlin et al., 2019) for serialization and then passed through a Transformer to obtain a latent representation $z^{l}$ with the same dimensionality as $z^{m}$.
@@ 3.2 Experts
To improve motion specialization and sim-to-real transfer, we introduce expert policies, i.e. motion tracking and delta action policies, trained on motion clusters derived from AE embeddings. All models in this part are implemented using a three-layer multilayer perceptron (MLP) and are trained via reinforcement learning (RL). More details of reward design can be found in the Appendix B.
@@ 3.2.1 Motion tracking training
Initially, a general motion tracking policy is trained on the entire dataset as a base model. Note that both joint angles and selected keypoint positions of the tracking/reference motion are included in the policy’s observation. More details about the observation design of our framework are included in the Appendix A.
Instead of training from scratch, all expert motion tracking policies are derived from the base model. This is because we expect each expert to retain some generalization capability for other motion clusters. Each expert policy is then fine-tuned on a specific motion cluster, allowing the policy to specialize in a behaviorally consistent cluster of skills, e.g., walking, standing, or jumping. This fine-tuning significantly improves motion fidelity and effectively resolves training conflicts caused by diverse motion types.
@@ 3.2.2 Multi-stage delta action training
In more detail, each expert motion tracking policy is deployed on the real robot to collect its corresponding motion trajectories in the real-world environment. At each timestep $t$, using the onboard sensors of the 29-DoF Unitree G1 robot (23 DoF actively controlled, excluding the wrist joints), we record the following: base linear velocity $v_{t}^{\text{base}}\in\mathbb{R}^{3}$, the robot base orientation as a quaternion $\alpha_{t}^{\text{base}}\in\mathbb{R}^{4}$, the base angular velocity $\omega_{t}^{\text{base}}\in\mathbb{R}^{3}$, the vector of joint positions $q_{t}\in\mathbb{R}^{23}$, and joint velocities $\dot{q}_{t}\in\mathbb{R}^{23}$. For each expert, we randomly sample several dozen reference motions for repeated execution, resulting in the collection of over a hundred real-world trajectories in total. Subsequently, we further train expert delta action models based on the data collected for each expert following ASAP (He et al., 2025). We find that, due to the consistent dynamic characteristics within each cluster of motion, this expert-specific training significantly improves the fitting accuracy of the delta action models and enables more efficient correction of the sim-to-real gap. Compared to a general delta action model, the expert models exhibit notable advantages in motion compensation accuracy and overall control performance.
@@ 3.3 Generalist
After optimizing the expert policies, we employ knowledge distillation to integrate the knowledge of each expert policy and generate a general whole-body control policy. We adopt DAgger (Ross et al., 2011) to implement multi-experts distillation. The distillation loss function is defined as:
@@
[…]
However, we observed that the capacity of the three-layer MLP is limited, making it insufficient to effectively learn behaviors of multiple expert policies. To address this issue, we adopt a more expressive architecture, the Transformer, to serve as the backbone for the final general policy model. The Transformer enables better modeling of complex patterns across diverse states, allowing for more effective fusion of expert knowledge. More details of the architecture can be found in the Appendix C.
@@ 4.1 Experiment Setup
OmniH2O (He et al., 2024a) 85.65% 87.83 0.2630 15.64% 360.96 0.4601
[…]
Hover (He et al., 2024c) 63.21% 105.84 0.2792 16.12% 323.08 0.3428
[…]
@@
To evaluate the effectiveness of our method, we compare it with three state-of-the-art (SOTA) methods: OmniH2O (He et al., 2024a), Exbody2 (Ji et al., 2024), and Hover (He et al., 2024c). For the fair comparison, we either use the officially released code or closely follow the official implementation when code is unavailable, adapting each method to the Unitree G1 robot (instead of H1/H1-2). In addition, we keep the training dataset consistent with that used for BB. For Hover, we use unmasked observations during the evaluation to ensure optimal performance.
@@
We assess performance using three key metrics: Success Rate (SR), Mean Per Joint Position Error (MPJPE), and Mean Per Keypoint Position Error (MPKPE). SR reflects the overall capability of the policy. MPJPE measures its accuracy in tracking retargeted joint angles, while MPKPE evaluates the precision of tracking retargeted keypoints in the world coordinate frame. Among all metrics, SR is the most critical, as it reflects the overall viability and stability of the policy. Other metrics, such as MPJPE and MPKPE, are only meaningful when the policy can successfully complete the task, as indicated by a high SR. Detailed definitions of these metrics can be found in Appendix A.
@@ 4.2 Cluster Analysis
[…]
How do we determine the number of clusters? We determine the number of clusters using the Elbow Method (Thorndike, 1953), which identifies the trade-off point between using fewer clusters and minimizing the within-cluster sum of squares (total square distance from all points to their respective cluster centers). Based on this analysis, we set the number of clusters to six, as shown in Figure 2.
@@ 4.4 Expert Performance
Refer to caption(a) Jump(b) Stand Low(c) Stand Up(d) Walk Slow
@@
Do expert policies outperform the generalist within their respective motion clusters? To answer this question, we evaluate the performance of expert policies. In Figure 3, we show performance improvements across four motion clusters during multi-stage delta action training, with results for the remaining clusters included in Appendix D. Specifically, we compare success rates in MuJoCo across four variants: the generalist policy without expert specialization (General Init), expert policies before delta action fine-tuning (Expert Init), expert policies after two fine-tuning rounds (Expert Final), and the final distilled generalist (General Final).
@@
We take stand-low as an example to provide a more intuitive understanding. As shown in Figure 4, without delta action fine-tuning (Iter 0), the policy fails to stabilize the foot, resulting in landing failure and deployment breakdown. After the first iteration (Iter 1), landing stability improves significantly, although the robot still struggles to lift its feet and exhibits visible trembling. After the second iteration (Iter 2), the policy can track the motion smoothly and maintain balance. More detailed visualizations can be found in the Appendix D.
@@
Why does a single delta action model struggle to generalize across motion clusters? To further investigate the variation in delta actions across different motion clusters, we analyze the magnitude of the delta action outputs. As shown in Table 6, Stand-up exhibits the smallest delta values, which is expected given that its motions are largely static. In contrast, Run shows the highest delta magnitudes due to its inherently large motion amplitudes. These distributional shifts across clusters introduce conflicts when attempting to train a single delta action model, further highlighting the need for cluster-specific modeling.

### v3

@@ Abstract
Achieving general agile whole-body control on humanoid robots remains a major challenge due to diverse motion demands and data conflicts. While existing frameworks excel in training single motion-specific policies, they struggle to generalize across highly varied behaviors due to conflicting control requirements and mismatched data distributions. In this work, we propose BumbleBee (BB), an expert-generalist learning framework that combines motion clustering and sim-to-real adaptation to overcome these challenges. BB first leverages an autoencoder-based clustering method to group behaviorally similar motions using motion features and motion descriptions. Expert policies are then trained within each cluster and refined with real-world data through iterative delta action modeling to bridge the sim-to-real gap. Finally, these experts are distilled into a unified generalist controller that preserves agility and robustness across all motion types. Experiments on two simulators and a real humanoid robot demonstrate that BB achieves state-of-the-art general whole-body control, setting a new benchmark for agile, robust, and generalizable humanoid control in the real world.
[…]
[Date]September 1, 2025
@@ 1 Introduction
Humanity has long dreamed of using humanoid robots to assist or replace humans in completing various daily tasks, including grabbing objects [1, 2, 3, 4, 5], carrying goods [6, 7, 8], etc. At present, a single agile whole-body humanoid skill can be trained to a remarkably high level of performance. The ASAP [9] framework first trains motion-specific tracking policies in simulation, then refines them using real-world data through motion-specific dynamic policy learning. This framework enables the robot to effectively bridge the sim-to-real gap and achieve a wide repertoire of agile and expressive behaviors. However, it is hard to directly apply this framework to general motion tracking, as different motions require the robot to focus on different aspects of control. For example, aggressive actions like jumping or fast walking demand precise, high-torque control, while conservative motions prioritize balance and smoothness. These differences lead to mismatched data distributions across motion types. It can lead to conflicting gradients during training, ultimately degrading the performance of the policy.
@@
Intuitively, to mitigate the issues caused by mismatched data distributions, there are two main directions. One is at the model level (Be more powerful): employing more expressive architectures such as Transformers or Diffusion Models [10, 11, 12, 13], which are capable of capturing complex motion distributions across diverse skills. The other is at the data level (Decompose the complexity): introducing structure into the training data, such as clustering motions by type or difficulty, balancing data across motion categories, or using curriculum learning strategies to gradually expose the policy to increasingly diverse behaviors [14]. These approaches aim to reduce interference between conflicting samples and improve the generalization capability of the learned policy. Inspired by the success of mixture of experts (MoE) [15, 16] in LLMs, we follow the second path to build a general whole-body control policy.
@@
To this end, we propose BumbleBee (BB), an expert-generalist pipeline that enables a general agile humanoid whole-body controller. Since both tracking and dynamic policies benefit from training on behaviorally similar motions, the dataset is first clustered using an autoencoder (AE) [17] trained on kinematic features, including additional leg-related features, and motion text descriptions. This design reflects the fact that different leg movements, such as walking, jumping, or crouching, require distinct torque control patterns, making leg dynamics a key factor in categorizing motion. A general tracking policy is initially trained on the full dataset and then fine-tuned within each cluster to obtain specialized expert policies. These experts are deployed on the real robot to collect real-world trajectories. For each cluster, a delta action model is trained to compensate for sim-to-real discrepancies by modeling the difference between simulated and real-world state transitions. The expert policies are further refined using the corresponding delta model in an iterative manner. Finally, a unified generalist is obtained through a combination of expert knowledge distillation.
@@ 2.1 Humanoid Whole-Body Control
Whole-body control (WBC) [18, 19, 20, 21, 22, 23, 24, 25] is essential for high DoF humanoid robots to perform coordinated and stable movements in complex tasks. Recent advances in reinforcement learning have enabled robots to acquire diverse whole-body skills, yet training a general WBC policy remains challenging due to hardware constraints and the inherent complexity of the robotic action space.
Previous works address this in different ways: the H2O series [26, 27] incorporates global keypoint positions but suffers from error accumulation in long-horizon tasks; HumanPlus [28] uses only joint angles for better real-world adaptability; Hover [29] introduces observation masking for flexible control mode switching. However, these methods are all trained on the full AMASS [30] dataset, where diverse motion distributions may cause gradient conflicts and reduce generalization. Exbody2 [14] attempts to mitigate this through progressive learning based on task difficulty.
@@ 2.2 Sim-to-Real Transfer
The sim-to-real transfer challenge primarily arises from two factors: (1) inaccuracies in robot modeling, including structural simplifications and parameter errors, and (2) the complex, nonlinear, and time-varying dynamics of real-world environments, e. g. contact events, friction variations, and elastic deformations, that exceed the fidelity of modern physics simulators.
Traditional approaches address this gap through System Identification (SysID) [31, 32, 33, 34, 35], which uses real-world observations to calibrate simulator parameters for improved physical realism. Domain Randomization offers an alternative by perturbing environment parameters during training to enhance policy robustness and generalization.
@@ 3 Method
3.1 Dataset Curation
3.1.1 Human Motion Retargeting
Following prior work [26], we first retarget SMPL-format [36] human motion sequences from the AMASS dataset into robot-specific representations, including global translations and joint axis-angle rotations. Given that AMASS contains a wide range of motions, such as crawling and climbing, we perform an additional data cleaning step to ensure quality. We apply PHC [37] to filter the dataset [9], resulting in 8,179 high-quality trajectories for clustering and training.
3.1.2 AE Clustering
@@
Our method not only relies solely on autoencoders to reconstruct full motion sequences, but also incorporates textual annotations to align motions at the semantic level. This ensures that motions with diverse patterns but shared semantics are positioned closer in the latent space. For example, the act of walking may take the form of linear movement or circular paths. These patterns would be hard to associate with based only on motion alignment. However, things can be different when textual semantic alignment is introduced. Note that the text information for training is from the open-source HumanML3D [38] dataset, which offers textual annotations and corresponding frame ranges for most of the sequences in AMASS.
@@
For the encoding process, we adopt a Transformer for motion encoding, following the previous work [39, 40]. The motion encoder takes as input a motion sequence $M_{\rm full}=\{\mathbf{p}_{t},\mathbf{r}_{t},\dot{\mathbf{r}}_{t},\mathbf{c}_{t},\mathbf{v}^{\text{feet}}_{t}\}_{t=1}^{T}$ and outputs a motion representation $z^{m}$, where $\mathbf{p}_{t}\in\mathbb{R}^{N\times 3}$ represents the 3D positions of all joints, $\mathbf{r}_{t}\in\mathbb{R}^{3}$ is the root translation in 3D space, $\dot{\mathbf{r}}_{t}\in\mathbb{R}^{3}$ is the root velocity, $\mathbf{c}_{t}\in\{0,1\}^{F}$ denotes the binary contact states of $F$ feet, and $\mathbf{v}^{\text{feet}}_{t}\in\mathbb{R}^{F\times 3}$ represents the 3D velocities of the $F$ feet. Textual data $l$ is first processed through a BERT model [41] for serialization and then passed through a Transformer to obtain a latent representation $z^{l}$ with the same dimensionality as $z^{m}$.
@@ 3.2 Experts
To improve motion specialization and sim-to-real transfer, we introduce expert policies, i. e. motion tracking and delta action policies, trained on motion clusters derived from AE embeddings. All models in this part are implemented using a three-layer multilayer perceptron (MLP) and are trained via reinforcement learning (RL). More details of reward design can be found in the Appendix 7.
@@ 3.2.1 Motion Tracking Training
Initially, a general motion tracking policy is trained on the entire dataset as a base model. Note that both joint angles and selected keypoint positions of the tracking/reference motion are included in the policy’s observation. More details about the observation design of our framework are included in the Appendix 6.
Instead of training from scratch, all expert motion tracking policies are derived from the base model. This is because we expect each expert to retain some generalization capability for other motion clusters. Each expert policy is then fine-tuned on a specific motion cluster, allowing the policy to specialize in a behaviorally consistent cluster of skills, e. g. walking, standing, or jumping. This fine-tuning significantly improves motion fidelity and effectively resolves training conflicts caused by diverse motion types.
@@ 3.2.2 Multi-Stage Delta Action Training
In more detail, each expert motion tracking policy is deployed on the real robot to collect its corresponding motion trajectories in the real-world environment. At each timestep $t$, using the onboard sensors of the 29-DoF Unitree G1 robot (23 DoF actively controlled, excluding the wrist joints), we record the following: base linear velocity $v_{t}^{\text{base}}\in\mathbb{R}^{3}$, base orientation as a quaternion $\alpha_{t}^{\text{base}}\in\mathbb{R}^{4}$, base angular velocity $\omega_{t}^{\text{base}}\in\mathbb{R}^{3}$, the vector of joint positions $q_{t}\in\mathbb{R}^{23}$, and joint velocities $\dot{q}_{t}\in\mathbb{R}^{23}$. For each expert, we randomly sample several dozen reference motions for repeated execution, resulting in the collection of over a hundred real-world trajectories in total. Subsequently, we further train expert delta action models based on the data collected for each expert following ASAP [9]. We find that, due to the consistent dynamic features within each cluster of motion, this expert-specific training significantly improves the fitting accuracy of the delta action models and enables more efficient correction of the sim-to-real gap. Compared to a general delta action model, the expert models exhibit notable advantages in motion compensation accuracy and overall control performance.
@@ 3.3 Generalist
After optimizing the expert policies, we employ knowledge distillation to integrate the knowledge of each expert policy and generate a general whole-body control policy. We adopt DAgger [42] to implement multi-experts distillation. The distillation loss function is defined as:
@@
where $\mathbb{E}_{s\sim\mathcal{D}}$ denotes the expectation over the states $s$ in the training dataset $\mathcal{D}$, $\text{KL}(\cdot\|\cdot)$ is the Kullback-Leibler divergence, $p_{\text{expert},k(s)}$ is the expert policy corresponding to state $s$, and $p_{\text{general}}$ is the general policy.
However, we observed that the capacity of the three-layer MLP is limited, making it insufficient to effectively learn behaviors of multiple expert policies. To address this issue, we adopt a more expressive architecture, the Transformer, to serve as the backbone for the final general policy model. The Transformer enables better modeling of complex patterns across diverse states, allowing for more effective fusion of expert knowledge. More details of the architecture can be found in the Appendix 8.
@@ 4.1 Experiment Setup
OmniH2O [26] 85.65% 87.83 0.2630 15.64% 360.96 0.4601
Exbody2 [14] 86.63% 86.66 0.2937 50.19% 272.42 0.3576
Hover [29] 63.21% 105.84 0.2792 16.12% 323.08 0.3428
BumbleBee (BB) 89.58% 83.30 0.1907 66.84% 294.27 0.2356
@@
To evaluate the effectiveness of our method, we compare it with three state-of-the-art (SOTA) methods: OmniH2O [26], Exbody2 [14], and Hover [29]. For the fair comparison, we either use the officially released code or closely follow the official implementation when code is unavailable, adapting each method to the Unitree G1 robot (instead of H1/H1-2). In addition, we keep the training dataset consistent with that used for BB. For Hover, we use unmasked observations during the evaluation to ensure optimal performance.
@@
We assess performance using three key metrics: Success Rate (SR), Mean Per Joint Position Error (MPJPE), and Mean Per Keypoint Position Error (MPKPE). SR reflects the overall capability of the policy. MPJPE measures its accuracy in tracking retargeted joint angles, while MPKPE evaluates the precision of tracking retargeted keypoints in the world coordinate frame. Among all metrics, SR is the most critical, as it reflects the overall viability and stability of the policy. Other metrics, such as MPJPE and MPKPE, are only meaningful when the policy can successfully complete the task, as indicated by a high SR. Detailed definitions of these metrics can be found in Appendix 6.
@@ 4.2 Cluster Analysis
Jump 2.32 9.21 0.329 0.379 0.346 0.405 jumps, jumping
Walk-slow 3.42 1.60 0.353 0.279 0.358 0.215 jogs, runs
Walk-fast 3.24 5.63 0.429 0.430 0.472 0.396 walks, forward
Stand-up 0.89 0.68 0.061 0.230 0.136 0.305 something, hand
Stand-mid 1.33 0.82 0.119 0.198 0.170 0.221 arms, hand
Stand-low 1.84 1.52 0.148 0.274 0.281 0.267 foot, leg
How do we determine the number of clusters? We determine the number of clusters using the Elbow Method [43], which identifies the trade-off point between using fewer clusters and minimizing the within-cluster sum of squares (total square distance from all points to their respective cluster centers). Based on this analysis, we set the number of clusters to six, as shown in Figure 2.
@@ 4.4 Expert Performance
Refer to caption(a) Stand Low(b) Jump(c) Stand Up(d) Walk Slow
@@
Do expert policies outperform the generalist within their respective motion clusters? To answer this question, we evaluate the performance of expert policies. In Figure 3, we show performance improvements across four motion clusters during multi-stage delta action training, with results for the remaining clusters included in Appendix 9. Specifically, we compare success rates in MuJoCo across four variants: the generalist policy without expert specialization (General Init), expert policies before delta action fine-tuning (Expert Init), expert policies after two fine-tuning rounds (Expert Final), and the final distilled generalist (General Final).
@@
We take stand-low as an example to provide a more intuitive understanding. As shown in Figure 4, without delta action fine-tuning (Iter 0), the policy fails to stabilize the foot, resulting in landing failure and deployment breakdown. After the first iteration (Iter 1), landing stability improves significantly, although the robot still struggles to lift its feet and exhibits visible trembling. After the second iteration (Iter 2), the policy can track the motion smoothly and maintain balance. More detailed visualizations can be found in the Appendix 9.
@@
Why does a single delta action model struggle to generalize across motion clusters? To further investigate the variation in delta actions across different motion clusters, we analyze the magnitude of the delta action outputs. As shown in Table 6, Stand-up exhibits the smallest delta values, which is expected given that its motions are largely static. In contrast, Walk-fast shows the highest delta magnitudes due to its inherently large motion amplitudes. These distributional shifts across clusters introduce conflicts when attempting to train a single delta action model, further highlighting the need for cluster-specific modeling.